Two tuning forks of frequencies 256 and 258 vibrations/second are sounded together. Then the time interval between two consecutive maxima heard by an observer is
(a) 2 sec
(b) 0.5 sec
(c) 250 sec
(d) 252 sec

Grade:12

2 Answers

Arun
3 years ago
n1=256n2=258Bothtuningforksaremakingbeatsso,weknowthatt(formax.)=n2n11sec=2582561=0.5sec
Hence, the option B is the correct answer.
Vikas TU
3 years ago
T = 1/n1-n2
= 1/258-256
= 1/2 ….............................................................................................................